# HGCLIP: Exploring Vision-Language Models with Graph Representations for Hierarchical Understanding

### v1

@@ Abstract
Object categories are typically organized into a multi-granularity taxonomic hierarchy. When classifying categories at different hierarchy levels, traditional uni-modal approaches focus primarily on image features, revealing limitations in complex scenarios. Recent studies integrating Vision-Language Models (VLMs) with class hierarchies have shown promise, yet they fall short of fully exploiting the hierarchical relationships. These efforts are constrained by their inability to perform effectively across varied granularity of categories. To tackle this issue, we propose a novel framework (HGCLIP) that effectively combines CLIP with a deeper exploitation of the Hierarchical class structure via Graph representation learning. We explore constructing the class hierarchy into a graph, with its nodes representing the textual or image features of each category. After passing through a graph encoder, the textual features incorporate hierarchical structure information, while the image features emphasize class-aware features derived from prototypes through the attention mechanism. Our approach demonstrates significant improvements on both generic and fine-grained visual recognition benchmarks. Our codes are fully available at https://github.com/richard-peng-xia/HGCLIP.
@@ 1 Introduction
Figure 1: An illustration of the graph representation based on class hierarchy. (a) The class hierarchy is presented in a tree structure. (b) The hierarchical structure is constructed into a graph, with nodes representing the textual/image features of each category. The graph is fed into a graph encoder, where the nodes update their parameters by aggregating the messages from their neighboring nodes. Thus, the class features are fused with hierarchical structural information via graph representation learning.
@@
Hierarchical image classification [38, 13] aims to enhance classification accuracy by identifying objects at various levels of granularity and capturing subtle relationships among them. Specifically, all the classes are organized into a multi-granularity taxonomic hierarchy (see in Figure 0(a)), where the top-level nodes represent broader categories (“Mammal”), while the lower-level nodes encompass finer-grained subcategories (“Dog”). The inherently hierarchical nature of the task compounds its complexity, as models must exhibit a keen understanding of semantic hierarchies, balancing the trade-off between capturing fine-grained details for subclasses while maintaining a broad understanding of superclasses [4]. Previous works [3, 13] mainly focus on enhancing image features according to the hierarchy of multiple branch outputs. These uni-modal methods only focus on the image modality, leading to certain limitations in complex scenarios, such as the inability to effectively utilize the textual descriptions of hierarchical labels and adapt to new classes or datasets. Therefore, leveraging multi-modal models (e.g., VLMs) to address hierarchical image classification presents stronger potential, offering richer information and greater scalability.
@@
Given the powerful generalization capabilities of VLMs [36, 19, 52] demonstrated on downstream tasks, harnessing their capabilities to address hierarchical image classification tasks presents a highly valuable exploration. These models are pre-trained on large-scale text-image pairs to align features from the image and text modalities in a shared latent embedding space. The predicted probabilities are obtained by calculating the similarity between image features and text features.
@@
Recently, some works have explored improving accuracy based on VLMs via class hierarchies. Specifically, CHiLS [32] employs hierarchical mappings to transform each class into a list of subcategories. However, this approach has significant drawbacks when applied to fine-grained datasets, as the subcategories of these labels tend to be specialized and rare, resulting in an overly detailed and contextually sparse representation. Utilizing these specific labels as prompts may overwhelm the model, lacking broader contextual relevance. Hierarchy-CLIP [11] proposes a label augmentation method that leverages the WordNet [9] label hierarchy, enriching each class with its parent and child classes. This method aims to provide a richer semantic expansion of class descriptions. It enhances surface-level semantic associations rather than delving into the deeper and more structured connections inherent in a hierarchical structure. This limitation becomes apparent in scenarios requiring classification across multiple hierarchical levels, where a nuanced understanding of these relationships is crucial. Moreover, these methods are both training-free. While this offers the advantage of simplicity and direct application, it lacks the capacity for further model adaptation to specific datasets. Additionally, these methods do not fully exploit the potential of VLMs to adapt to the diverse and complex nature of hierarchical understanding.
@@
To address this issue, we first introduce prompt learning [55, 54, 22] as an efficient method to adapt VLMs to downstream tasks. HGCLIP introduces prompt tokens within the multi-modal branches of CLIP to facilitate the learning of hierarchical contextual representations. More importantly, as demonstrated in Figure 0(b), HGCLIP explores the integration of CLIP with graph representations for hierarchical image classification. Specifically, hierarchical relationships are modeled as a graph, given that they inherently form a tree-like structure. Based on this graph, we employ a graph encoder [24, 15, 43] to encode text features, enabling them to incorporate hierarchical structural information. Moreover, since image features represent features of individual patches/pixels rather than categories, we utilize prototype learning to represent image features for each category. Similarly, a graph encoder is leveraged to allow the prototypes to learn hierarchical relationships, and subsequently utilize the attention mechanism to enable the spatial feature map of images to focus more on the class-aware features derived from prototypes. On hierarchical image classification, HGCLIP outperforms existing CLIP-based approaches across both generic and fine-grained datasets. In scenarios where hierarchical labels are unavailable, HGCLIP also improves accuracy when utilizing class hierarchies queried by ChatGPT [33]. Further, HGCLIP demonstrates favorable generalization ability and robustness in domain generalization and subpopulation shift settings, resulting in consistent improvements over existing methods. To sum up, the main contributions of this work include:
  • We propose HGCLIP, a SoTA method in hierarchical image classification for adaptation of CLIP.
@@
  • Our approach exhibits new state-of-the-art (SoTA) performance across eight hierarchical image classification benchmarks and performs commendably on eight extra datasets with distribution shifts.
@@ 2 Related Work
Prompt Learning in Vision-Language Models: VLMs leverage information from both image and text modalities to encode multimodal representations. VLMs, e.g., CLIP [36], ALIGN [19], and LiT [52] are pre-trained on large-scale image-text pairs and have demonstrated remarkable representation abilities on various downstream tasks [10, 57, 7]. However, efficiently adapting them to downstream tasks is still a major challenge. Prompt learning [28, 27], as a parameter-efficient technique, is well-suited for utilizing the representation capacity of pre-trained VLMs to boost performance, instead of the resource-intensive process of full fine-tuning. Many works [55, 22, 46, 40, 56, 14] have demonstrated powerful performance on specific downstream tasks by combining VLMs and prompt tuning.
@@
Hierarchical Image Classification: Hierarchical image classification [38, 13] aims to categorize images into a hierarchical structure that organizes classes into a tree-like taxonomy. It acknowledges the inherent hierarchical nature of visual concepts, allowing for more nuanced and contextually rich image categorization. Prior research has explored various methodologies, including model architectures tailored for hierarchical classification [13, 3, 4], and exploiting the relationship of the categories in the hierarchy [29]. Furthermore, the development of hierarchical classification has spurred some works [21, 49] that harness the class hierarchy across diverse domains, as these models tend to focus more on fine-grained and semantically relevant features. Recently, hierarchical labels are integrated with VLMs [32, 11]. Nonetheless, these methods roughly overlook the hierarchical relationships among labels. Our work comprehensively leverages the hierarchical relationships among labels, resulting in performance improvements on both generic and fine-grained datasets.
@@
Graph Representation Learning: Modern graph analysis methods rely on graph representation learning, encompassing graph embedding, graph neural networks (GNNs), and transformers. Early graph embedding techniques [35, 41, 12] typically map nodes into a low-dimensional space, capturing structural information like proximity between nodes [2]. Recently, GNNs [24, 15, 42, 47] have become the mainstream technique in graph representation learning. They rely on a message-passing framework where each node refines its representation by recursively aggregating messages from its neighbors [45]. Moreover, some recent approaches have also explored transformer-based architectures [51, 18, 50]. Hierarchical graph representations are employed in this work to enrich multi-modal features, thus improving the model performance and generalization.
@@ 3 Preliminaries
In this work, our goal is to learn hierarchical textual and visual knowledge via graph encoder based on CLIP. We introduce related concepts and definitions in the following.
@@ 3.1 Revisiting CLIP
We denote the CLIP image and text encoder as $\mathcal{I}(\cdot)$ and $\mathcal{T}(\cdot)$. The dataset contains $K$ categories, i.e., $\{C_{1},\cdots,C_{K}\}$. CLIP leverages a structured approach by inserting all category names into a predefined textual template represented by the [CLASS] token, e.g., creating expressions like “a photo of a [CLASS].”. This results in the generation of textual inputs denoted as $T_{K}$. Subsequently, textual features, represented as $\mathbf{F}_{t}\in\mathbb{R}^{K\times D}$, are extracted. Each input image $I$ is divided into $M$ fixed-sized patches, and each patch is embedded into $D$-dimensional latent space. Then CLIP derives its spatial feature map $\mathbf{F}_{s}\in\mathbb{R}^{H\times W\times D}$ and computes the global visual representations $\mathbf{f}_{v}\in\mathbb{R}^{1\times D}$ through pooling operations, where $H$ and $W$ denote the height and width of the feature map. The integration of features from both encoders is achieved through cosine similarity measures, ultimately yielding classification $logits\in\mathbb{R}^{1\times K}$. This comprehensive process can be summarized as follows
@@
[…]
Figure 3: t-SNE plots of image embeddings in previous SoTA prompting method CoCoOp, MaPLe, and HGCLIP on two datasets with distinct semantic granularities. HGCLIP shows better separability in both fine-grained and coarse-grained levels.
@@ 3.2 Graph Encoder
Graph. A graph is represented as $G=(V,E)$, with $V$ denoting the set of nodes and $E$ the set of edges. Equivalently, the graph can be represented by an adjacency matrix $A$, such as $A_{ij}=1$, if $(v_{i},v_{j})\in E$, for any $v_{i},v_{j}\in V$. Graph Encoder. GNNs are popular choices of graph encoder, most of which employ a message-passing mechanism [45]. Specifically, each node in the graph aggregates messages (i.e., input features or embeddings) from its neighboring nodes to update its own embedding. Multiple layers of neighborhood aggregation can be stacked, facilitating recursive message passing across the graph. Formally, in the $l$-th GNN layer, the embedding of node $v$, denoted by $\mathbf{f}^{l}_{v}$, is calculated based on the embeddings in the previous layer, as follows
@@
where $\mathcal{N}_{v}$ is the set of neighboring nodes of $v$, $\theta^{l}$ is the learnable GNN parameters in layer $l$. $\textsc{Aggr}(\cdot)$ is the neighborhood aggregation function and can take various forms, ranging from the simple mean pooling [24, 15] to advanced neural networks such as neural attention [42] or multi-layer perceptrons [47]. Note that the initial node embedding $\mathbf{f}^{0}_{v}$ is simply given by the input feature. We abstract the multi-layer encoding process as
@@
where $\Theta=(\theta^{1},\ldots,\theta^{L})$ is the collection of weights across the layers. Note that graph embedding methods [35, 41, 12] and graph transformers [51, 18, 50] could also serves as GraphEncoder.
@@ 4.1 Hierarchy Setting
The ground truth class hierarchy currently available in a dataset is usually obtained by querying a WordNet [9]-like dictionary, but in the real world, our dataset may have no available class hierarchy. In this case, we turn to LLMs, i.e., ChatGPT [33] (GPT-3.5-Turbo), to approximate the hierarchy diagram. Specifically, given some label set size $K$, semantic granularity levels $h$, class names, and optional context, we query ChatGPT with the prompt:
@@ 4.2 Multi-modal Hierarchical Prompt
In order to comprehensively and efficiently leverage the capabilities of pretrained VLMs, we explore the potential of multi-modal prompt, encompassing both textual and visual prompt. As highlighted in [22], the acquisition of prompt at deeper transformer layers is crucial, as it progressively models hierarchical feature representations. Learnable tokens are introduced at multiple transformer blocks of both textual and visual branches of VLMs, given as textual prompt $\mathbf{P}^{T}=\{\mathbf{p}^{T}_{1},\cdots,\mathbf{p}^{T}_{t}\}$ and visual prompt $\mathbf{P}^{V}=\{\mathbf{p}^{V}_{1},\cdots,\mathbf{p}^{V}_{v}\}$, respectively. Therefore, the image encoder processes the input tokens added visual prompt $\mathbf{P}^{V}$ to generate prompted spatial feature map represented as $\tilde{\mathbf{F}_{s}}\in\mathbb{R}^{(HW+v)\times D}$ and prompted global visual representations $\tilde{\mathbf{f}_{v}}\in\mathbb{R}^{1\times D}$. Similarly, textual prompt $\mathbf{P}^{T}$ are incorporated into the input tokens for encoding, and textual features are obtained as $\tilde{\mathbf{F}_{t}}\in\mathbb{R}^{K\times D}$. These hierarchical prompt tokens leverage the knowledge encoding capabilities of VLMs to effectively learn task-relevant contextual representations across different semantic levels.

### v3

@@ Abstract
Object categories are typically organized into a multi-granularity taxonomic hierarchy. When classifying categories at different hierarchy levels, traditional uni-modal approaches focus primarily on image features, revealing limitations in complex scenarios. Recent studies integrating Vision-Language Models (VLMs) with class hierarchies have shown promise, yet they fall short of fully exploiting the hierarchical relationships. These efforts are constrained by their inability to perform effectively across varied granularity of categories. To tackle this issue, we propose a novel framework (HGCLIP) that effectively combines CLIP with a deeper exploitation of the Hierarchical class structure via Graph representation learning. We explore constructing the class hierarchy into a graph, with its nodes representing the textual or image features of each category. After passing through a graph encoder, the textual features incorporate hierarchical structure information, while the image features emphasize class-aware features derived from prototypes through the attention mechanism. Our approach demonstrates significant improvements on 11 diverse visual recognition benchmarks. Our codes are fully available at https://github.com/richard-peng-xia/HGCLIP.
HGCLIP: Exploring Vision-Language Models with Graph Representations for Hierarchical Understanding
Peng Xia1,4, Xingtong Yu3, Ming Hu1, Lie Ju1, Zhiyong Wang2, Peibo Duan1, Zongyuan Ge1 1Monash University, 2 The University of Sydney, 3 Singapore Management University, 4 UNC-Chapel Hill pxia@cs.unc.edu, zongyuan.ge@monash.edu
@@ 1 Introduction
Hierarchical image classification Salakhutdinov et al. (2011); Guo et al. (2018) aims to enhance classification accuracy by identifying objects at various levels of granularity and capturing subtle relationships among them. Specifically, all the classes are organized into a multi-granularity taxonomic hierarchy (see in Figure 1a), where the top-level nodes represent broader categories (“Mammal”), while the lower-level nodes encompass finer-grained subcategories (“Dog”). The inherently hierarchical nature of the task compounds its complexity, as models must exhibit a keen understanding of semantic hierarchies, balancing the trade-off between capturing fine-grained details for subclasses while maintaining a broad understanding of superclasses Chen et al. (2018). Previous works Chang et al. (2021); Guo et al. (2018) mainly focus on enhancing image features according to the hierarchy of multiple branch outputs. These uni-modal methods only focus on the image modality, leading to certain limitations in complex scenarios, such as the inability to effectively utilize the textual descriptions of hierarchical labels and adapt to new classes or datasets. Therefore, leveraging multi-modal models (e.g., VLMs) to address hierarchical image classification presents stronger potential, offering richer information and greater scalability.
Figure 1: An illustration of the graph representation based on class hierarchy. (a) The class hierarchy is presented in a tree structure. (b) The hierarchical labels are constructed into a graph, with nodes representing the text/image features of each class. The graph is fed into a graph encoder, where the nodes update the parameters by aggregating the messages from their neighboring nodes. Thus, the class features are fused with hierarchical information via graph representation learning.
@@
Given the powerful generalization capabilities of VLMs Radford et al. (2021); Jia et al. (2021); Zhai et al. (2022) demonstrated on downstream tasks, harnessing their capabilities to address hierarchical image classification tasks presents a highly valuable exploration. These models are pre-trained on large-scale text-image pairs to align features from the image and text modalities in a shared latent embedding space. The predicted probabilities are obtained by calculating the similarity between image features and text features.
@@
Recently, some works have explored improving accuracy based on VLMs via class hierarchies. Specifically, CHiLS Novack et al. (2023) employs hierarchical mappings to transform each class into a list of subcategories. However, this approach has significant drawbacks when applied to fine-grained datasets, as the subcategories of these labels tend to be specialized and rare, resulting in an overly detailed and contextually sparse representation. Utilizing these specific labels as prompts may overwhelm the model, lacking broader contextual relevance. Hierarchy-CLIP Ge et al. (2023) proposes a label augmentation method that leverages the WordNet Fellbaum (2010) label hierarchy, enriching each class with its parent and child classes. This method aims to provide a richer semantic expansion of class descriptions. It enhances surface-level semantic associations rather than delving into the deeper and more structured connections inherent in a hierarchical structure. This limitation becomes apparent in scenarios requiring classification across multiple hierarchical levels, where a nuanced understanding of these relationships is crucial. Moreover, these methods are both training-free. While this offers the advantage of simplicity and direct application, it lacks the capacity for further model adaptation to specific datasets. Additionally, these methods do not fully exploit the potential of VLMs to adapt to the diverse and complex nature of hierarchical understanding.
@@
To address this issue, we first introduce prompt learning Zhou et al. (2022b); Khattak et al. (2023a) as an efficient method to adapt VLMs to downstream tasks. HGCLIP introduces prompt tokens within the multi-modal branches of CLIP to facilitate the learning of hierarchical contextual representations. More importantly, as demonstrated in Figure 1b, HGCLIP explores the integration of CLIP with graph representations for hierarchical image classification. Specifically, hierarchical relationships are modeled as a graph, given that they inherently form a tree-like structure. Based on this graph, we employ a graph encoder Veličković et al. (2018a) to encode text features, enabling them to incorporate hierarchical structural information. Moreover, since image features represent features of individual patches/pixels rather than categories, we utilize prototype learning to represent image features for each category. Similarly, a graph encoder is leveraged to allow the prototypes to learn hierarchical relationships, and subsequently utilize the attention mechanism to enable the spatial feature map of images to focus more on the class-aware features derived from prototypes. On hierarchical image classification, HGCLIP outperforms existing CLIP-based approaches across both generic and fine-grained datasets. In scenarios where hierarchical labels are unavailable, HGCLIP also improves accuracy when utilizing class hierarchies queried by ChatGPT OpenAI (2023). Further, HGCLIP demonstrates favorable generalization ability and robustness in domain generalization and subpopulation shift settings, resulting in consistent improvements over existing methods. To sum up, the main contributions of this work include:
  • We propose HGCLIP, a state-of-the-art (SoTA) method in hierarchical image classification for adaptation of CLIP.
@@
  • Our approach exhibits new SoTA performance across eleven hierarchical image classification benchmarks.
@@ 2 Related Work
Prompt Learning in Vision-Language Models: VLMs leverage information from both image and text modalities to encode multimodal representations. VLMs, e.g., CLIP Radford et al. (2021), ALIGN Jia et al. (2021), and LiT Zhai et al. (2022) are pre-trained on large-scale image-text pairs and demonstrate remarkable representation abilities on various downstream tasks Gao et al. (2023); Zhu et al. (2022); Ding et al. (2022); Xia et al. (2024c, e, d, a). However, efficiently adapting them to downstream tasks is still a major challenge. Prompt learning Li and Liang (2021); Lester et al. (2021), as a parameter-efficient technique, is well-suited for utilizing the representation capacity of pre-trained VLMs to boost performance, instead of the resource-intensive process of full fine-tuning. Many works Shu et al. (2022); Zhou et al. (2023); Li et al. (2024) have demonstrated powerful performance on specific downstream tasks by combining VLMs and prompt tuning.
@@
Hierarchical Image Classification: Hierarchical image classification Salakhutdinov et al. (2011); Guo et al. (2018) aims to categorize images into a hierarchical structure that organizes classes into a tree-like taxonomy. It acknowledges the inherent hierarchical nature of visual concepts, allowing for more nuanced and contextually rich image categorization. Prior research has explored various methodologies, including model architectures tailored for hierarchical classification Guo et al. (2018); Chang et al. (2021); Chen et al. (2018), and exploiting the relationship of the categories in the hierarchy Liu et al. (2022). Furthermore, the development of hierarchical classification has spurred some works Ju et al. (2023, 2024); Yi et al. (2022) that harness the class hierarchy across diverse domains, as these models tend to focus more on fine-grained and semantically relevant features. Recently, hierarchical labels are integrated with VLMs Novack et al. (2023); Ge et al. (2023). Nonetheless, these methods roughly overlook the hierarchical relationships among labels. Our work comprehensively leverages the hierarchical relationships among labels, resulting in performance improvements on both generic and fine-grained datasets.
@@
Graph Representation Learning: Modern graph analysis methods rely on graph representation learning, encompassing graph embedding, graph neural networks (GNNs), and transformers. Early graph embedding techniques Perozzi et al. (2014); Grover and Leskovec (2016) typically map nodes into a low-dimensional space, capturing structural information. Recently, GNNs Kipf and Welling (2016); Veličković et al. (2018a) have become the mainstream technique in graph representation learning. They rely on a message-passing framework where each node refines its representation by recursively aggregating messages from its neighbors. Moreover, some recent approaches have also explored transformer-based architectures Yun et al. (2019); Hu et al. (2020). Furthermore, the boom of graph representation learning also advances the research and development in other communities such as CV Shi et al. (2019) and NLP Zhang et al. (2023). In this work, we employ hierarchical graph representations to enrich multi-modal features, thus improving the model performance and generalization.
Figure 2: t-SNE plots of image embeddings in SOTA method CoCoOp, MaPLe, and HGCLIP on two datasets with distinct semantic granularities. HGCLIP shows better separability in both fine-grained and coarse-grained levels.
@@ 3 Preliminaries
In this work, our goal is to learn hierarchical multi-modal knowledge via graph encoder based on CLIP. We will introduce related concepts and definitions in the following.
@@ 3.1 Revisiting CLIP
We denote the CLIP image and text encoder as $\mathcal{I}(\cdot)$ and $\mathcal{T}(\cdot)$. The dataset contains $K$ categories, i.e., $\{C_{1},\cdots,C_{K}\}$. CLIP leverages a structured approach by inserting all category names into a predefined textual template represented by the [CLASS] token, e.g., creating expressions like “a photo of a [CLASS].". This results in the generation of textual inputs denoted as $T_{K}$. Subsequently, textual features, represented as $\mathbf{F}_{t}\in\mathbb{R}^{K\times D}$, are extracted. Each input image $I$ is divided into $M$ fixed-sized patches, and each patch is embedded into $D$-dimensional latent space. Then CLIP derives its spatial feature map $\mathbf{F}_{s}\in\mathbb{R}^{H\times W\times D}$ and computes the global visual representations $\mathbf{f}_{v}\in\mathbb{R}^{1\times D}$ through pooling operations, where $H$ and $W$ denote the height and width of the feature map. The integration of features from both encoders is achieved through cosine similarity measures, ultimately yielding classification $logits\in\mathbb{R}^{1\times K}$. This comprehensive process can be summarized as follows
@@
$\textit{logits}=\mathbf{f}_{v}{\mathbf{F}_{t}}^{T}.$ (3)
The matrix multiplication operation between $\mathbf{f}_{v}$ and $\mathbf{F}_{t}$ is equivalent to calculating cosine similarities, assumed to be $L_{2}$-normalized features. logits signifies the computed probabilities for all $K$ categories, and CLIP identifies the category with the maximum output probability $\mathrm{argmax}_{C_{K}}(\textit{logits})$ as its final prediction.
@@ 3.2 Graph Encoder
Graph. A graph is represented as $G=(V,E)$, with $V$ denoting the set of nodes and $E$ the set of edges. Equivalently, the graph can be represented by an adjacency matrix $A$, such as $A_{ij}=1$, if $(v_{i},v_{j})\in E$, for any $v_{i},v_{j}\in V$.
Graph Encoder. GNNs are popular choices of graph encoder, most of which employ a message-passing mechanism Wu et al. (2020). Specifically, each node in the graph aggregates messages (i.e., input features or embeddings) from its neighboring nodes to update its own embedding. Multiple layers of neighborhood aggregation can be stacked, facilitating recursive message passing across the graph. Formally, in the $l$-th GNN layer, the embedding of node $v$, denoted by $\mathbf{f}^{l}_{v}$, is calculated based on the embeddings in the previous layer, as follows
@@
where $\mathcal{N}_{v}$ is the set of neighboring nodes of $v$, $\theta^{l}$ is the learnable GNN parameters in layer $l$. $\textsc{Aggr}(\cdot)$ is the neighborhood aggregation function and can take various forms, ranging from the simple mean pooling Kipf and Welling (2016) to advanced neural networks such as neural attention Veličković et al. (2018a) or multi-layer perceptrons Xu et al. (2019). Note that the initial node embedding $\mathbf{f}^{0}_{v}$ is simply given by the input feature. We abstract the multi-layer encoding process as
@@
where $\Theta=(\theta^{1},\ldots,\theta^{L})$ is the collection of weights across the layers. Note that graph embedding methods Perozzi et al. (2014); Tang et al. (2015); Grover and Leskovec (2016) and graph transformers Yun et al. (2019); Hu et al. (2020); Ying et al. (2021) could also serves as GraphEncoder.
@@ 4.1 Hierarchy Setting
The ground truth class hierarchy currently available in a dataset is usually obtained by querying a WordNet Fellbaum (2010)-like dictionary, but in the real world, our dataset may have no available class hierarchy. In this case, we turn to LLMs, i.e., ChatGPT, to approximate the hierarchy diagram. Specifically, given some label set size $K$, semantic granularity levels $h$, class names, and optional context, we query ChatGPT with the prompt:
@@ 4.2 Multi-modal Hierarchical Prompt
In order to comprehensively and efficiently leverage the capabilities of pretrained VLMs, we explore the potential of multi-modal prompt, encompassing both textual and visual prompt. As highlighted in Khattak et al. (2023a), the acquisition of prompt at deeper transformer layers is crucial, as it progressively models hierarchical feature representations. Learnable tokens are introduced at multiple transformer blocks of both textual and visual branches of VLMs, given as textual prompt $\mathbf{P}^{T}=\{\mathbf{p}^{T}_{1},\cdots,\mathbf{p}^{T}_{t}\}$ and visual prompt $\mathbf{P}^{V}=\{\mathbf{p}^{V}_{1},\cdots,\mathbf{p}^{V}_{v}\}$, respectively. Therefore, the image encoder processes the input tokens added visual prompt $\mathbf{P}^{V}$ to generate prompted spatial feature map represented as $\tilde{\mathbf{F}_{s}}\in\mathbb{R}^{(HW+v)\times D}$ and prompted global visual representations $\tilde{\mathbf{f}_{v}}\in\mathbb{R}^{1\times D}$. Similarly, textual prompt $\mathbf{P}^{T}$ are incorporated into the input tokens for encoding, and textual features are obtained as $\tilde{\mathbf{F}_{t}}\in\mathbb{R}^{K\times D}$. These hierarchical prompt tokens leverage the knowledge encoding capabilities of VLMs to effectively learn task-relevant contextual representations across different semantic levels.